mathslover (mathslover):

An African elephant can reach heights of 13 feet and possess a mass of as much as 6000 kg. Determine the weight of an African elephant in Newtons and in pounds. (Given: 1.00 N = .225 pounds)

OpenStudy (anonymous):

W(N)=60000N W(P)=60000*.225=13500

mathslover (mathslover):

@zepp @apoorvk @Diyadiya

OpenStudy (zepp):

6000kg = ? pounds?

mathslover (mathslover):

1kg = 2.20 pounds 6000 kg = 6000*2.2 pounds = 13200

mathslover (mathslover):

@keshavsarda please explain

OpenStudy (anonymous):

W=mg

OpenStudy (zepp):

Okay good, give that 1 newton is equal to 0.225 pounds \(\large \frac{1N}{0.225lbs}=\frac{xN}{13200}\) Solve for x :)

mathslover (mathslover):

oh thanska lot @zepp and @keshavsarda

OpenStudy (zepp):

yw :)

mathslover (mathslover):

i have a quiz tomorrow and i am preparing for it can any1 just give me some important introduction to newtons's laws of motion @JamesJ can u help me ?

OpenStudy (jamesj):

Your request is very broad. My best recommendation would be to watch the first few videos of this course: http://ocw.mit.edu/courses/physics/8-01-physics-i-classical-mechanics-fall-1999/video-lectures/ Lecture 6 in particular if you want to focus on the laws of motion.

mathslover (mathslover):

Oh ! thanks a lot for ur comment sir ... i will surely focus on that video .. and i hope that i will be able to do my best for my team and for myself